# Peekaboo: Interactive Video Generation via Masked-Diffusion

### 12 Dec 2023 v1

@@ Abstract
Recently there has been a lot of progress in text-to-video generation, with state-of-the-art models being capable of generating high quality, realistic videos. However, these models lack the capability for users to interactively control and generate videos, which can potentially unlock new areas of application. As a first step towards this goal, we tackle the problem of endowing diffusion-based video generation models with interactive spatio-temporal control over their output. To this end, we take inspiration from the recent advances in segmentation literature to propose a novel spatio-temporal masked attention module - Peekaboo. This module is a training-free, no-inference-overhead addition to off-the-shelf video generation models which enables spatio-temporal control. We also propose an evaluation benchmark for the interactive video generation task. Through extensive qualitative and quantitative evaluation, we establish that Peekaboo enables control video generation and even obtains a gain of upto 3.8×3.8$\times$3.8 × in mIoU over baseline models.
Figure 1: Zero-training interactive video generation. Peekaboo allows users to control the output (object size, location and motion) for off-the-shelf video diffusion models, through specially designed masking modules. First row shows a panda playing Peekaboo by following an expanding mask on left.
@@ 1 Introduction
Generating realistic videos from natural language descriptions is a challenging but exciting task that has recently made significant progress  [32, 18, 37, 35]. This is largely due to the development of powerful generative models and latent diffusion models (LDMs [29]), which can produce high-quality and diverse videos from text. These models have opened up new possibilities for creative applications and expression.
@@ 2.1 Video Generation
Text-based video generation using latent diffusion model has taken a significant leap in recent years [32, 14, 41, 13, 7]. Make-a-video [32] introduced the 3D UNet architecture, by decomposing attention layers into spatial, cross and temporal attention layers. Further progress in this generation pipeline was made by [15, 41, 13, 7], while keeping the core three attention-layer architecture intact. Although these works focus on generating videos with high relevance to the text input, they do not provide spatio-temporal control in each frame. More recent works have tried to equip models with this ability to control generation spatio-temporally. Such methods have integrated guidance from depth maps [11], target motion [6, 16] or a combination of these modalities to generate videos [36]. However, all these works either require re-training the base model or an external adapter with aligned grounded spatio-temporal data, which is a challenging and expensive task.
@@ 4.3 Zero-training Pipeline
Putting the selective masks in a diffusion pipeline gives us a zero-training method, dubbed Peekaboo. Peekaboo integrates in the attention layers of the 3D-UNet architecture of text-to-video models. We perform selective generation of foreground and background object for a fixed number of steps $t$ and then allow free-generation for the rest of steps. This free generation enables the foreground and background pixels to cohesively integrate with each other on the same canvas as have been done by [21, 3]. In essence, our method ensures that foreground pixels cannot “see” the background pixels for some steps (and vice versa), before being visible to each other. This is akin to a game of Peekaboo.
@@ 4.4 Extensions
Automatically generated input masks
Since our method is orthogonal to the choice of input masks, we can use a large language model to generate the input masks for an object corresponding to a given prompt, in a similar fashion as concurrent works [23, 21]. In Table 2, we demonstrate that doing this leads to videos with better quality than the baseline model. Moreover, it enables our method to be end-to-end in terms of only requiring a text prompt from the user.
@@
Image generation
Image generation diffusion method are based of 2D-UNet architecture, with the absence of temporal attention layer. Analogous to our text-to-video setup, we can adapt Peekaboo for Image Diffusion models. The spatial-attention mask maintains the semantic structure of the image while the cross-attention mask focuses the attention of foreground token on desired location and vice versa for background. In Figure 7, we showcase spatial control on an off-the-shelf diffusion model and highlight the versatility of our method.
@@ 5.1.1 Spatial Control
Method Peekaboo ssv2-ST Interactive Motion Control (IMC)
mIoU % ($\uparrow$) Coverage % ($\uparrow$) CD ($\downarrow$) AP50 % ($\uparrow$) mIoU % ($\uparrow$) Coverage % ($\uparrow$) CD ($\downarrow$) AP50 % ($\uparrow$)
ZeroScope [35] - 13.9 42.0 0.22 9.3 12.6 88 0.26 0.6
34.7 56.3 0.17 39.8 36.3 96.3 0.12 33.8
ModelScope [35] - 12.0 44.7 0.17 6.6 9.6 93.3 0.25 2.35
33.2 63.7 0.10 35.8 36.1 96.6 0.13 33.3
@@
Table 1: Evaluation of spatio-temporal control: We evaluate two different models for video generation with spatio-temporal control on ssv2-ST and IMC datasets. As demonstrated by mIoU and CD, the videos generated by Peekaboo endow the baselines with spatio-temporal control. Peekaboo also increases the quality of the main objects in the scene, as seen by higher coverage and AP50 scores.
Evaluation Datasets
@@
Evaluating spatial control in multiple text-to-video models is a challenging task and requires creating a common benchmark for (prompt, mask) pairs. We develop a benchmark obtained from a public video dataset with high-quality masks that represent realistic locations for day-to-day subjects. Further, we also curated a set of (prompt, mask) pairs that represent an interactive input from the user in controlling a video and its subject.
@@
  • Something-something v2-Spatio-Temporal (ssv2-ST): We use Something-Something v2 dataset [12, 24] to obtain the generation prompts and ground truth masks from real action videos. We filter out a set of 295 prompts. The details for this filtering are in the appendix. We then use an off-the-shelf OWL-ViT-large open-vocabulary object detector [25] to obtain the bounding box annotations of the object in the videos. This set represents bounding box and prompt pairs of real-world videos, serving as a test bed for both the quality and control of methods for generating realistic videos with spatio-temporal control.
@@
  • Interactive Motion Control (IMC): We also curate a set of prompts and bounding boxes which are manually defined. We use GPT-4 to generate prompts and pick a set of 34 prompts of objects in their natural contexts. These prompts are varied in the type of object, size of the object and the type of motion exhibited. We then annotate 3 sets of bounding boxes for each prompt, where the location, path taken, speed and size are varied. This set of 102 prompt-bounding box pairs serve as our custom evaluation set for spatial control. Note that since ssv2-ST dataset has a lot of inanimate objects, we bias this dataset to contain more living objects. This dataset represents possible input pairs that real users may generate.
@@
We use two base models for our evaluation, Zeroscope and ModelScope [35]. These models are run for the default number of inference steps, with default temperature and classifier guidance parameters. We also experiment with mask guidance steps in the appendix. We provide the model with the text prompt and the set of input bounding boxes. The generated videos are then evaluated for spatio-temporal control and video quality.
Evaluation methodology.
After generating videos for each (prompt, mask) pair, we pass these videos through OWL-ViT-large detector to compute bounding boxes for each generated video. We first compute the fraction of generated videos for which OwL-ViT detects bounding boxes in more that 50% of the generated frames. We report this fraction as the Coverage of the model in Table 1. However, the lack of a detected bounding box does not necessarily imply the lack of an object generated, since OwL-ViT could fail to capture some objects correctly. Hence, to evaluate the spatio-temporal control of the generation method, we first filter out videos where less than 50% frames have a detected bounding box. We then compute the Intersection-over-Union of the detected bounding boxes and the input mask on these filtered videos. We report the mean of these IoU (mIoU) scores for each method in Table 1. These two metrics together provide a good proxy of the quality of the generated videos as well as the spatio-temporal control imparted. We compute the Centroid Distance (CD) as the distance between the centroid of the generated object and input mask, normalized to 1. This measures control of the generation location. Finally, we report the average precision@50% (AP50) of the detected and input bounding boxes averaged over all videos. For generated frames with the object present, AP50 represents the spatial control provided by the method, while mIoU measures the model’s ability to match the input bounding boxes exactly and penalizes frames where the object cannot be detected.
Figure 4: (a) AP50 Scores for our datasets: Subplot (a) shows the performance of baseline models with and without Peekaboo on IMC and ssv2-ST. Our method provides a large gain in the scores. (b) Ablation Studies on IMC: The performance of Peekaboo varies as different attention masks are removed. The AP50 drops the most when cross-attention masks are removed, indicating their importance to spatial control, followed by temporal and spatial attention. Best viewed in color.
@@
Results
In Table 1, we demonstrate that our method adds control to the model. We verify that our method enables spatio-temporal control, as evidenced by the lower CD and higher (upto 2.5x) AP@50 scores on both the IMC and ssv2-ST dataset. This means that the generated objects are close to the true centroid of the input mask, and their shape and size are also consistent with the input mask. We observe significant jump in mIoU score with Peekaboo across different models, highlighting superior spatio-temporal control achieved through Peekaboo. Finally, we note that Peekaboo has a higher coverage than the baseline models, indicating that our method is also able to generate objects when the base model could not do so.
@@ 5.1.2 Quality control
While the above datasets provide evidence for Peekaboo’s spatio-temporal control, we also benchmark our method on MSR-VTT[39] – a large scale video generation dataset – to evaluate the quality of videos generated. We benchmark Peekaboo for evaluating quality control using Fréchet Video Distance score (FVD) metric [34]. FVD is calculated based on I3D model trained on Kinetics-400 dataset [5]. Following previous works, we evaluate on the test-set of MSR-VTT containing 2900 videos by randomly sampling one of the 20 captions for each video. We demonstrate the versatility of our method by using bounding boxes generated by GPT-4. We query GPT-4 to generate series of locations for the foreground object depending on the prompt. We evaluate on ModelScope model and compare the scores with Peekaboo. Table 2 shows that Peekaboo increases the quality of generated while providing spatial control during video generation. The performance of these methods is also better than other baselines, indicating that Peekaboo can be integrated in an automated pipeline to use GPT-4 generated bounding boxes and output a coherent video.
@@
(c) A spider descending on its web.
(f) A croissant on a wooden table.
Figure 6: Overcoming model failures: Frames on the left are generated by zero-scope, and frames on the right are generated by Peekaboo. Inset in the first row are cross-attention map between the word “spider” and the pixels in the video frame. We can generate objects that are otherwise omitted from the video by the base model. The attention maps also show that explicit masking leads to better generation. The second row depicts a numeracy failure of the baseline where Peekaboo can control the number of objects.
@@ 5.1.3 Ablation analysis
A Spatio-Temporal attention block consists of three types of attention layers– Spatial, Cross and Temporal. Peekaboo applies masking on all three layers, however, the effect of each mask on the generation quality is different. In this section, we experiment with Peekaboo by disabling masking for each attention layer one-by-one. We evaluate the AP50 score for ModelScope and ZeroScope on the IMC dataset, as shown in Figure 4. The performance drops massively when any one of the attention mask is not provided. We observe that not passing $M_{CA}$ hurts the control the most. This is explained by the fact that main object’s text token will not focus its attention at the bounding box location, leading to the object being generated at a different location. Surprisingly, not passing $M_{TA}$ is worse than not passing $M_{SA}$. We conjecture that removing spatial attention mask leads to degraded videos, while removing the temporal attention mask leads to the loss of temporal control. Since the latter model still generates higher quality objects at incorrect locations, it has a lower AP50 score. We notice that the Coverage of the model after removing $M_{SA}$ is much less than the Coverage of the model after removing $M_{TA}$, providing evidence in support of our hypothesis.
@@ 5.2 Qualitative Results
In Figure 1, we present examples of videos generated by Peekaboo applied on ZeroScope [35]. As demonstrated, the videos follow the bounding box input. Through these qualitative results, we highlight the versatility of bounding box input in capturing the shape, size, location and motion, and show how our method can utilize this information interactively.
Static spatial control.
Figure 3 shows videos where the object is statically located in the frame. Our method can control the position of the object, and can also change the size of the object as specified by the user through a bounding box.
Dynamic spatial control.
Figure 5 present videos where the main subject is moving on a desired path. Our method generated realistic looking movements for various motion trajectories. The temporal masking of our method also enables it to handle cases where the mask disappears mid-way through the scene, as is the case in the first row in Figure 1, while the spatial and cross-attention masking ensures spatial coherence of the generated frames with the input bounding boxes.
Overcoming model failures.
Diffusion models can have a bias on their generation capabilities depending on their training data. However, we observe that Peekaboo can suppress those biases and produce high quality generation by forcing the model to generate foreground object at a specific location. In Figure 6, we present results of prompts where the original model fails to produce the foreground object however, our method can produce the object in the user specified location and motion. The inset figures in Figure 6 reveal the reason for this – while the cross-attention corresponding to the word “spider” is diffused across the entire canvas in the original model, Peekaboo focuses this attention on the desired region. Further, Figure 6 depicts the example of hallucination by generation model where the subject was generated multiple times. Again, Peekaboo solves this issue due to spatial-attention mask and cross-attention at a specific location.

### 19 Apr 2024 v2

@@ Abstract
Modern video generation models like Sora have achieved remarkable success in producing high-quality videos. However, a significant limitation is their inability to offer interactive control to users, a feature that promises to open up unprecedented applications and creativity. In this work, we introduce the first solution to equip diffusion-based video generation models with spatio-temporal control. We present Peekaboo, a novel masked attention module, which seamlessly integrates with current video generation models offering control without the need for additional training or inference overhead. To facilitate future research, we also introduce a comprehensive benchmark for interactive video generation. This benchmark offers a standardized framework for the community to assess the efficacy of emerging interactive video generation models. Our extensive qualitative and quantitative assessments reveal that Peekaboo achieves up to a 3.8×3.8$\times$3.8 × improvement in mIoU over baseline models, all while maintaining the same latency. Code and benchmark are available on the webpage.
Figure 1: Zero-training No-latency interactive video generation. Peekaboo allows users to control the output (object size, location and trajectory) for any off-the-shelf video diffusion models, through specially designed masking modules. First row shows a panda playing Peekaboo by following an expanding mask in left direction.
$\dagger$\$$\dagger$\$footnotetext: equal contribution
@@ 1 Introduction
Generating realistic videos from natural language descriptions is a challenging but exciting task that has recently made significant progress  [4, 35, 20, 40, 38]. This is largely due to the development of powerful generative models and latent diffusion models (LDMs [32]), which can produce high-quality and diverse videos from text. These models have opened up new possibilities for creative applications and expression.
@@ 2.1 Video Generation
Text-based video generation using latent diffusion model has taken a significant leap in recent years [4, 35, 16, 44, 15, 8]. Make-a-video [35] introduced the 3D UNet architecture, by decomposing attention layers into spatial, cross and temporal attention layers. Further progress in this generation pipeline was made by [17, 44, 15, 8], while keeping the core three attention-layer architecture intact. Although these works focus on generating videos with high relevance to the text input, they do not provide spatio-temporal control in each frame. More recent works have tried to equip models with this ability to control generation spatio-temporally. Such methods have integrated guidance from depth maps [12], target motion [7, 18] or a combination of these modalities to generate videos [39]. However, all these works either require re-training the base model or an external adapter with aligned grounded spatio-temporal data, which is a challenging and expensive task.
@@ 4.3 Zero-training Pipeline
Putting the selective masks in a diffusion pipeline gives us a zero-training method, dubbed Peekaboo. Peekaboo integrates in the attention layers of the 3D-UNet architecture of text-to-video models. We perform selective generation of foreground and background object for a fixed number of steps $t$ and then allow free-generation for the rest of steps. This free generation enables the foreground and background pixels to cohesively integrate with each other on the same canvas as have been done by [23, 3]. In essence, our method ensures that foreground pixels cannot “see" the background pixels for some steps (and vice versa), before being visible to each other. This is akin to a game of Peekaboo.
@@ 4.4 Extensions
Automatically generated input masks Since our method is orthogonal to the choice of input masks, we can use a large language model to generate the input masks for an object corresponding to a given prompt, in a similar fashion as concurrent works [25, 23]. In Table 2, we demonstrate that doing this leads to videos with better quality than the baseline model. Moreover, it enables our method to be end-to-end in terms of only requiring a text prompt from the user.
@@
Image generation Image generation diffusion method are based of 2D-UNet architecture, with the absence of temporal attention layer. Analogous to our text-to-video setup, we can adapt Peekaboo for Image Diffusion models. The spatial-attention mask maintains the semantic structure of the image while the cross-attention mask focuses the attention of foreground token on desired location and vice versa for background. In Figure 7, we showcase spatial control on an off-the-shelf diffusion model and highlight the versatility of our method.
@@ 5.1.1 Spatial Control
Method Latency DAVIS16 LaSOT ssv2-ST IMC
mIoU ($\uparrow$) AP50 ($\uparrow$) Cov ($\uparrow$) CD ($\downarrow$) mIoU AP50 Cov CD ($\downarrow$) mIoU AP50 Cov CD ($\downarrow$) mIoU AP50 Cov CD ($\downarrow$)
LLM-VD [24]* 2.20$\times$ 26.1 15.2 96 0.19 13.5 4.6 98 0.24 27.2 21.2 61 0.12 36.1 33.3 97 0.13
ModelScope [38] 1$\times$ 19.6 5.7 100 0.25 4.0 0.7 96 0.33 12.0 6.6 44.7 0.17 9.6 2.4 93.3 0.25
w/ Peekaboo 1.03$\times$ 26.0 16.6 93 0.18 14.6 10.2 98 0.25 33.2 35.8 63.7 0.10 36.1 33.3 96.6 0.13
ZeroScope [38] 1$\times$ 11.7 0.1 100 0.22 3.6 0.4 100 0.3 13.9 9.3 42.0 0.22 12.6 0.6 88.0 0.26
w/ Peekaboo 1.03$\times$ 20.6 17.9 100 0.19 11.5 11.9 100 0.28 34.7 39.8 56.3 0.17 36.3 33.8 96.3 0.12
@@
Table 1: Evaluation of spatio-temporal control on mIoU, AP50, Coverage and Centroid Distance (CD): We evaluate two different video generation models on spatio-temporal control against DAVIS16, LaSOT, ssv2-ST, and IMC datasets. As demonstrated by mIoU and CD, the videos generated by Peekaboo endow the baselines with spatio-temporal control. Peekaboo also increases the quality of the main objects in the scene, as seen by higher AP50 and Coverage scores. Further, LLM-VD[24] has higher inference cost whereas Peekaboo does not affect latency. *: LLM-VD[24] has not released code, this is our re-implementation.
@@
Evaluation Datasets Evaluating spatial control in multiple text-to-video models is a challenging task and requires creating a common benchmark for (prompt, mask) pairs. We develop a benchmark obtained from a public video dataset with high-quality masks that represent realistic locations for day-to-day subjects. Further, we also curated a set of (prompt, mask) pairs that represent an interactive input from the user in controlling a video and its subject.
@@
  • Something-something v2-Spatio-Temporal (ssv2-ST): We use Something-Something v2 dataset [14, 26] to obtain the generation prompts and ground truth masks from real action videos. We filter out a set of 295 prompts. The details for this filtering are in the appendix. We then use an off-the-shelf OWL-ViT-large open-vocabulary object detector [27] to obtain the bounding box (bbox) annotations of the object in the videos. This set represents bbox and prompt pairs of real-world videos, serving as a test bed for both the quality and control of methods for generating realistic videos with spatio-temporal control.
@@
  • Interactive Motion Control (IMC): We also curate a set of prompts and bounding boxes which are manually defined. We use GPT-4 to generate prompts and pick a set of 34 prompts of objects in their natural contexts. These prompts are varied in the type of object, size of the object and the type of motion exhibited. We then annotate 3 sets of bboxes for each prompt, where the location, path taken, speed and size are varied. This set of 102 prompt-bbox pairs serve as our custom evaluation set for spatial control. Note that since ssv2-ST dataset has a lot of inanimate objects, we bias this dataset to contain more living objects. This dataset represents possible input pairs that real users may generate.
  • LaSOT: We repurpose a large-scale object tracking dataset –LaSOT[13]– for evaluating control in video generation. This dataset contains prompt-bbox-video triplets for a large number of classes. The videos have frame level annotations specifying the location of the object in the video. We subsample the videos to 8 FPS and then randomly pick up 2 clips per video from the test set of this dataset. This gives us 450 total clips across 70 different object categories.
  • DAVIS-16: DAVIS-16[30] is another video object segmentation dataset that we consider. We take videos from its test set, manually annotating them with prompts. We use the provided segmentation masks to create input bboxes. This gives us 40 prompt-bbox pairs in total, where each video has a different subject.
@@
We use two base models for our evaluation, Zeroscope and ModelScope [38]. These models are run for the default number of inference steps, with default temperature and classifier guidance parameters. We also experiment with mask guidance steps in the appendix. We provide the model with the text prompt and the set of input bboxes. The generated videos are then evaluated for spatio-temporal control and video quality.
Evaluation methodology. After generating videos for each (prompt, mask) pair, we pass these videos through OWL-ViT-large detector to compute bboxes for each generated video. We first compute the fraction of generated videos for which OwL-ViT detects bboxes in more that 50% of the generated frames. We report this fraction as the Coverage of the model in Table 1. However, the lack of a detected bbox does not necessarily imply the lack of an object generated, since OwL-ViT could fail to capture some objects correctly. Hence, to evaluate the spatio-temporal control of the generation method, we first filter out videos where less than 50% frames have a detected bbox. We then compute the Intersection-over-Union of the detected bboxes and the input mask on these filtered videos. We report the mean of these IoU (mIoU) scores for each method in Table 1. These two metrics together provide a good proxy of the quality of the generated videos as well as the spatio-temporal control imparted. We compute the Centroid Distance (CD) as the distance between the centroid of the generated object and input mask, normalized to 1. This measures control of the generation location. Finally, we report the average precision@50% (AP50) of the detected and input bboxes averaged over all videos. For generated frames with the object present, AP50 represents the spatial control provided by the method, while mIoU measures the model’s ability to match the input bboxes exactly and penalizes frames where the object cannot be detected.
Figure 4: (a) Ablation Studies on IMC: The performance of Peekaboo varies as different attention masks are removed. The AP50 drops the most when cross-attention masks are removed, indicating their importance to spatial control, followed by temporal and spatial attention. (b) Performance against varying bbox sizes: mIoU score compared against varying bounding box sizes across time in videos. We observe that Peekaboo provides better control independent of bbox size variation. Best viewed in color.
@@
Results In Table 1, we demonstrate that our method adds control to the model. We verify that our method enables spatio-temporal control, as evidenced by the higher (upto 2.5x) AP@50 scores and lower CF on all four datasets (DAVIS16, LaSOT, ssv2-ST and IMC). This means that the generated objects are close to the true centroid of the input mask, and their shape and size are also consistent with the input mask. We observe significant jump in mIoU score with Peekaboo across different models and LLM-VD [24], highlighting superior spatio-temporal control achieved through Peekaboo. Further, Peekaboo has a higher coverage than the baseline models and LLM-VD, indicating that our method is also able to generate objects when the base model could not do so. Finally, we note that Peekaboo introduce minimal latency increase to the original method compared to LLM-VD’s 2.20$\times$ increase in inference time on ModelScope.
@@ 5.1.2 Quality control
While the above datasets provide evidence for Peekaboo’s spatio-temporal control, we also benchmark our method on MSR-VTT[42] – a large scale video generation dataset – to evaluate the quality of videos generated. We benchmark Peekaboo for evaluating quality control using Fréchet Video Distance score (FVD) metric [37]. FVD is calculated based on I3D model trained on Kinetics-400 dataset [6]. Following previous works, we evaluate on the test-set of MSR-VTT containing 2900 videos by randomly sampling one of the 20 captions for each video. We demonstrate the versatility of our method by using bounding boxes generated by GPT-4. We query GPT-4 to generate series of locations for the foreground object depending on the prompt. We evaluate on ModelScope model and compare the scores with Peekaboo. Table 2 shows that Peekaboo increases the quality of generated while providing spatial control during video generation. The performance of these methods is also better than other baselines, indicating that Peekaboo can be integrated in an automated pipeline to use GPT-4 generated bboxes and output a coherent video.
@@
(a) A spider descending on its web.
(b) A croissant on a wooden table.
Figure 6: Overcoming model failures: Frames on the left are generated by zero-scope, and frames on the right are generated by Peekaboo. Inset in the first row are cross-attention map between the word “spider" and the pixels in the video frame. We can generate objects that are otherwise omitted from the video by the base model. The attention maps also show that explicit masking leads to better generation. The second row depicts a numeracy failure of the baseline where Peekaboo can control the number of objects.
@@ 5.1.3 Ablation analysis
Ablation on individual attention masks. A Spatio-Temporal attention block consists of three types of attention layers– Spatial, Cross and Temporal. Peekaboo applies masking on all three layers, however, the effect of each mask on the generation quality is different. In this section, we experiment with Peekaboo by disabling masking for each attention layer one-by-one. We evaluate the AP50 score for ModelScope and ZeroScope on the IMC dataset, as shown in Figure 4. The performance drops massively when any one of the attention mask is not provided. We observe that not passing $M_{CA}$ hurts the control the most. This is explained by the fact that main object’s text token will not focus its attention at the bbox location, leading to the object being generated at a different location. Surprisingly, not passing $M_{TA}$ is worse than not passing $M_{SA}$. We conjecture that removing spatial attention mask leads to degraded videos, while removing the temporal attention mask leads to the loss of temporal control. Since the latter model still generates higher quality objects at incorrect locations, it has a lower AP50 score. We notice that the Coverage of the model after removing $M_{SA}$ is much less than the Coverage of the model after removing $M_{TA}$, providing evidence in support of our hypothesis.
Performance on varying Bbox sizes. Peekaboo can accommodate varying bbox sizes within videos. Fig 3 shows qualitative control on input bboxes with the output. Moreover, SSv2-ST dataset contain examples with varying bbox sizes. Thus, we further quantify our SSv2-ST results in Fig 4(b) by plotting mIoU scores against varying box sizes across time in videos. For each prompt of SSv2-ST, we compute the standard deviation of the bbox size (relative to the mean input) to highlight the variation across time. We observe a consistent improvement by Peekaboo on base model independent of box size variation.
@@ 5.2 Qualitative Results
In Figure 1, we present examples of videos generated by Peekaboo applied on ZeroScope [38]. As demonstrated, the videos follow the bbox input. Through these qualitative results, we highlight the versatility of bbox input in capturing the shape, size, location and motion, and show how our method can utilize this information interactively.
Static spatial control. Figure 3 shows videos where the object is statically located in the frame. Our method can control the position of the object, and can also change the size of the object as specified by the user through a bbox.
Dynamic spatial control. Figure 5 present videos where the main subject is moving on a desired path. Our method generated realistic looking movements for various motion trajectories. The temporal masking of our method also enables it to handle cases where the mask disappears mid-way through the scene, as is the case in the first row in Figure 1, while the spatial and cross-attention masking ensures spatial coherence of the generated frames with the input bounding boxes.
Overcoming model failures. Diffusion models can have a bias on their generation capabilities depending on their training data. However, we observe that Peekaboo can suppress those biases and produce high quality generation by forcing the model to generate foreground object at a specific location. In Figure 6, we present results of prompts where the original model fails to produce the foreground object however, our method can produce the object in the user specified location and motion. The inset figures in Figure 6 reveal the reason for this – while the cross-attention corresponding to the word “spider" is diffused across the entire canvas in the original model, Peekaboo focuses this attention on the desired region. Further, Figure 6 depicts the example of hallucination by generation model where the subject was generated multiple times. Again, Peekaboo solves this issue due to spatial-attention mask and cross-attention at a specific location.
6 Conclusion
Appendix A Text to image synthesis